40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 23 17:54:51 2014, Harry, Update, General, Coupling Improvements plus PER Measurement Setup PERSetup.png
    Reply  Thu Jul 24 17:37:19 2014, Harry, Update, General, Coupling Improvements plus PER Measurement Setup 
Message ID: 10271     Entry time: Thu Jul 24 17:37:19 2014     In reply to: 10264
Author: Harry 
Type: Update 
Category: General 
Subject: Coupling Improvements plus PER Measurement Setup 

Quote:

 Purpose

We wanted to improve the coupling into the fibers, because it's very rarely good enough to take measurements with, as the beam is obscured by random noise.

Additionally, we want to add some things to the current setup in order to better measure Polarization Extinction Ratio.

What Was Done

After flailing for several hours, Koji helped me couple the NPRO light into the fiber, using the fiber illuminator for alignment. The coupled optical power immediately jumped from 0-1uW to 5.6mW (around 11% coupling).

Q and I discussed the setup for measuring PER. In addition to the current setup, we added a half wave plate to control the angle of the polarization, in addition to the existing quarter wave plate, which corrects the beam for ellipticity.

PERSetup.png

Once everything was coupled, I started minimizing S-Polarization coming out of the first polarizing beam splitter, and maximizing the P-Polarization entering the fibers.

I did this by first varying the Quarter Wave plate to eliminate as much S Polarization as possible, and then, maintaining a constant differential in angle between QWP and HWP, I rotated them both to maximize power coupled into the fibers.

I measured 0.2 mW of S-Polarization, and 54.3 mW of P-Polarization.

At this point, a locking effort started, and I had to leave the 40m.

Moving Forward

 

Tomorrow, I would like to finish the setup of the PER measurement design. That is to say, add a collimator to the other end of the fiber, and align it with the second PBS.

And, of course, take a measurement of the Polarization Extinction Ratio of the fiber.

To eventually be implemented in Frequency Offset Locking. 

 

Today, I encountered a problem with the stage that holds the coupler, in that its ability to rotate unchecked causes coupling to degrade over time due to torsion in the fibers. Our solution was to stress-relieve the fiber with a clamp.

Unfortunately, this also meant losing coupling completely. It was re-coupled at up 72% efficiency. (Subsequent changes in the setup have decreased that to ~24%)

When I took preliminary measurements of the PER, it was significant, which was unexpected. Upon further discussion with Q, we concluded that since the fiber's fast axis hadn't been aligned with the light's polarization, I was getting multiple polarizations out the end of the fiber.

Subsequent measurements of the power contained in the two polarizations of the output light gave about 0.8% S-Polarization introduced by the fiber.

Tomorrow

I would like to find another collimator holder, to hold the output side of the fiber.

Also, I will spend more time aligning the fiber axes, and the second PBS in order to get a better (read: more reasonable) measurement of PER.

ELOG V3.1.3-